One year ago, you invested $2,980. Today it is worth $3,800.50. What rate of interest did you earn?

a. 27.26 percent

b. 27.53 percent

c. 21.59 percent

d. 22.02 percent

e. 27.48 percent
Answer #1

Let Rate of interest be r

Value one year ago = P = 2980

Value now = F = 3800.50

=> F = P(1+r)

=> 3800.50 = 2980(1+r)

=> r = (3800.50/2980) -1 = 0.2753 or 27.53 percent

Hence, option (b)